calculate the heat change at 100°C for each of the following and indicate whether heat was absorbed or released: kilojoules to vaporize 5.00 kg if water
1
Expert's answer
2018-02-13T06:28:41-0500
To determine the amount of heat needed to evaporate a certain mass of a liquid, its specific heat of vaporization must be multiplied by mass : Q = r·m.

Specific heat of vapor formation determines the amount of heat needed to evaporate 1 kg of liquid at a certain temperature (r, J/kg).

The internal energy of 1 kg of water vapor at 100 °C is greater than the internal energy of water at that temperature of 2257 kJ.

r = 2.3·106 J/kg

Q = 2.3·106 J/kg · 5 kg = 11.5·106 J = 11.5·103 kJ

During vaporisation 5 kg of water is absorbed 11.5·103 kJ of heat.